hi, i also need help when a question says give a slope of a line perpendicular to the given line. and they say y=-2/3x-4. i don't know how to do that.
Answer by nerdybill(6962)

When you multiply the slopes of two lines perpendicular lines, the result is -1.
(this is the "negative reciprocal")
.
From:
y=-2/3x-4
since it is in "slope-intercept" form we immediately see that the slope is:
-2/3
So, a line perpendicular must have a slope (m) such that:
(-2/3)m = -1
(2/3)m = 1
(2)m = 3
m = 3/2
.
So, the answer is 3/2